OpenStudy (anonymous):

272 is what percent of 680?

OpenStudy (anonymous):

It is 40 percent

OpenStudy (anonymous):

idk how i got it but its the answer.

OpenStudy (anonymous):

you got it right but I want to know how to do it

OpenStudy (anonymous):

\[272/680*100\]

OpenStudy (anonymous):

oh yes i remeber now the solution is \[\frac{ a }{ b } \times 100 = \]

OpenStudy (anonymous):

You beat me to it lol xD

OpenStudy (anonymous):

take it easy bro. ^_^

OpenStudy (anonymous):

Always think of it as PART over WHOLE times 100 = % Think of pizza. If you ate 3 pieces of a 12 piece pizza, what percentage of the pizza did you eat? Part over whole times 100. 3/12 x 100 = 25%